# The Peak Frequency and Luminosity of Synchrotron Emitting Shocks: from Non-Relativistic to Ultra-Relativistic Explosions

### v1

@@ 6.1 Energy
Finally, we discuss the results of Barniol Duran et al. (2013) in light of our present work. Barniol Duran et al. (2013) generalized Chevalier (1998) and similar equipartition methods (e.g., Scott & Readhead, 1977) to the case of ultra-relativistic sources. In many ways this is similar to the objective of our present work. However, there are significant differences between the two analyses. First, Barniol Duran et al. (2013) assume a purely power-law electron distribution and do not incorporate the effects of thermal electrons. As we have shown here, thermal electrons are critically important in ultra-relativistic shocks as they—and not the power-law electrons—determine the properties at the peak of the SED (see Figure 1 panel c). Second, Barniol Duran et al. (2013) express their results in terms of a dimensionless parameter $\eta\equiv\min(\nu_{m}/\nu_{\rm a},1)$ that is not specified in that work. Here $\nu_{m}$ is the characteristic (minimum) frequency of power-law electrons, and is analogous to $\nu_{\Theta}$ in the thermal electron scenario. Using the results of Barniol Duran et al. (2013) with $\eta$ treated as a constant $\sim\mathcal{O}(1)$ is therefore equivalent to the assumption that the peak of the SED occurs at $\nu_{\rm pk}\sim\nu_{\rm a}$. We have shown here instead that emission at the peak of the SED is typically optically-thin for ultra-relativistic shocks. To the extent that this conclusion applies also in the purely power-law electron scenario, then one would generally expect $\nu_{\rm a}\ll\nu_{\rm pk}\simeq\nu_{\rm m}$, or $\eta\gg 1$. The fact that $\eta\nsim 1$, and that the value of $\eta$ is not explicitly determined (it will, among other things, depend on $\Gamma$) therefore makes it difficult to compare the Barniol Duran et al. (2013) expressions with our present results.
@@ 7 Summary
  • For shocks governed by SSA: thermal electrons dominate the emissivity and absorption at frequency $\nu_{\rm pk}$ when $L_{\nu_{\rm pk}}>L_{\rm th}$ (Equation 8; solid grey curve in Figure 2). For luminous events that fall within this regime, the standard Chevalier (1998) relations (4.1,4.1) are no longer applicable. Instead, the physical properties of the shock should be estimated from Equations (4.2,4.2; see also Appendix A) or using the numerical procedure outlined in Appendix B (implemented in code available at https://github.com/bmargalit/thermal-synchrotron-v2).
@@
  • There is a maximum allowed peak luminosity $L_{\nu_{\rm pk}}$, such that $L_{\nu_{\rm pk}}\leq L_{\rm crit}$ (Equation C7; see also Equation 5.1 and Figure 2). This is a consequence of the transition between optically-thin and optically-thick emission at the peak of the SED, and is a generic feature of synchrotron shock models.777 The expression for $L_{\rm crit}$ is determined by properties of the thermal electron population (§5; Appendix C). However, a qualitatively similar upper limit would apply also in the (unphysical) case where thermal electrons are neglected. This gives $L_{\rm crit}\sim 2\times 10^{33}\,{\rm erg\,s}^{-1}\,{\rm Hz}^{-1}\,\epsilon_{e,-2}^{-2}\epsilon_{B,-1}^{-23/55}(\nu_{\rm pk}\bar{t}/5{\rm GHz}\times 100{\rm d})^{133/55}$. While the specific luminosity at low frequencies $\nu\ll\nu_{\rm pk}$ can exceed this limit (that is, $L_{\nu}>L_{\rm crit}$ is possible) when peak emission is optically-thin, the peak luminosity should always satisfy $L_{\nu_{\rm pk}}\leq L_{\rm crit}$. It is interesting to note from Figure 2 that no observed event falls above this limit (red curve), despite clear observational biases towards detecting more luminous events. We argue that the limit $L_{\rm crit}$ that we have derived here provides a natural explanation to this fact.
@@
The code developed in this analysis is publicly available at https://github.com/bmargalit/thermal-synchrotron-v2, and may be particularly useful when the analytic approximations derived in the main text are not sufficiently accurate. This code can be imported as a python module and used to infer physical properties of observed transients, for example by fitting model-generated flux densities to real data. It may also be useful for generating plots similar to Figures (2,5).
B.M. thanks Nayana A. J. for sharing a compilation of $\nu_{\rm pk}$ and $L_{\nu_{\rm pk}}$ for SNe which provided a basis for the sample of events shown in Figures 2,5. This research benefited from interactions that were funded by the Gordon and Betty Moore Foundation through Grant GBMF5076. (Harris et al., 2020), scipy (Virtanen et al., 2020), matplotlib (Hunter, 2007).
@@ Appendix B General Solution
are functions of the shock velocity and effective mass-loss rate. In the above, $f(\Theta)$ and $g(\Theta)$ (not to be confused with the volume filling factor $f$) are functions of the normalized electron temperature and are given by Equations (5,8) of Margalit & Quataert (2021).888 Note that there was a typo in the expression for $g(\Theta)$ presented in Equation (8) of Margalit & Quataert (2021). The correct expression is $g(\Theta)=\frac{(p-1)\gamma_{m}}{(p-1)(\gamma_{m}-1)-p+2}\left(\frac{\gamma_{m}}{3\Theta}\right)^{p-1}$. The electron temperature is itself a function of the shock velocity, and is determined by Equation (2) of Margalit & Quataert (2021) with $\Theta_{0}=\epsilon_{T}(\mu m_{p}/\mu_{e}m_{e})(\Gamma-1)$. In the limit $\Theta\gtrsim 1$ which is relevant for high-velocity shocks, $f(\Theta)\approx g(\Theta)\approx 1$. The coefficients $C_{j}(p)$ and $C_{\alpha}(p)$ are given by Equations (15,17) of Margalit & Quataert (2021). For $p=3$, these attain the values $C_{j}=2/3$ and $C_{\alpha}\simeq 28.6$. Finally, the function $I^{\prime}(x)$ is given by Equation (13) of Margalit & Quataert (2021), and is derived in Mahadevan et al. (1996).

### v2

@@ 6.1 Energy
Finally, we discuss the results of Barniol Duran et al. (2013) in light of our present work. Barniol Duran et al. (2013) generalized Chevalier (1998) and similar equipartition methods (e.g., Scott & Readhead, 1977) to the case of ultra-relativistic sources. In many ways this is similar to the objective of our present work. However, there are significant differences between the two analyses. First, Barniol Duran et al. (2013) assume a purely power-law electron distribution and do not incorporate the effects of thermal electrons. As we have shown here, thermal electrons are critically important in relativistic shocks as they—and not the power-law electrons—determine the properties at the peak of the SED (see Figure 1 panel c). Second, Barniol Duran et al. (2013) express their results in terms of a dimensionless parameter $\eta\equiv\min(\nu_{m}/\nu_{\rm a},1)$ that is not specified in that work. Here $\nu_{m}$ is the characteristic (minimum) frequency of power-law electrons, and is analogous to $\nu_{\Theta}$ in the thermal electron scenario. Using the results of Barniol Duran et al. (2013) with $\eta$ treated as a constant $\sim\mathcal{O}(1)$ is therefore equivalent to the assumption that the peak of the SED occurs at $\nu_{\rm pk}\sim\nu_{\rm a}$. We have shown here instead that emission at the peak of the SED is typically optically-thin for ultra-relativistic shocks. To the extent that this conclusion applies also in the purely power-law electron scenario then one would generally expect $\nu_{\rm a}\ll\nu_{\rm pk}\simeq\nu_{\rm m}$, or $\eta\gg 1$ (examination of Table 3 in Granot & Sari 2002 indicates that this is typically true for GRBs, for example). The fact that $\eta\nsim 1$, and that the value of $\eta$ is not explicitly determined (it will, among other things, depend on $\Gamma$) therefore makes it difficult to compare the Barniol Duran et al. (2013) expressions with our present results.
6.2 Microphysical Parameters
The results of §6.1 apply in the scenario where $\epsilon_{e}\ll\epsilon_{T}\lesssim 1$ (that is, $\delta\ll 1$). In the following subsection we discuss this assumption, which we consider to be physically well-motivated, albeit with significant uncertainties. Before doing so, we note that a minimum on the explosion energy is often estimated by instead adopting ‘equipartition’ microphysical parameters, i.e. $\epsilon_{e}\sim\epsilon_{B}\sim\mathcal{O}(1)$. One can immediately see from Equation (16) that such values of $\epsilon_{e}$ and $\epsilon_{B}$ indeed minimize the inferred energy. In this scenario $\epsilon_{e}\sim\epsilon_{T}$ (i.e., $\delta\sim 1$), so that $L_{\rm th}$ is very large and there is little to no phase space in which thermal electrons dominate peak emission. That is, if power-law electrons are assumed to be in equipartition with magnetic fields and thermal electrons + ions, then $L_{\rm th}\sim L_{\rm crit}$ such that $L_{\nu_{\rm pk}}$ is almost always smaller than $L_{\rm th}$, and therefore Equation (16)—and not (17)—is appropriate for estimating the energy. In other words, in the equipartition scenario thermal electrons are almost never relevant, even if they are present. This implies that the standard minimum-energy estimates, which neglect thermal electrons entirely, are still valid lower-limits on $U$.
The reduction in inferred energies when including thermal electrons that is shown in Figure 5 and discussed in §6.1 applies to the more physically-motivated case where $\epsilon_{e}\ll 1$ rather than equipartition. This would alleviate energetic constraints for more reasonable expected values of $\epsilon_{e}$ and $\epsilon_{B}$. We also note that an equipartition argument holds true also in the case where thermal electrons are included and dominant over power-law electrons ($\epsilon_{e}\ll\epsilon_{T}$, $L_{\nu_{\rm pk}}>L_{\rm th}$). This can be derived from Equation (17), which again shows that $U$ is minimized when $\epsilon_{T}\sim\epsilon_{B}\sim\mathcal{O}(1)$.
The details and efficiency of both non-thermal electron acceleration and electron-ion equilibration in collisionless shocks are in general not fully understood. This introduces significant uncertainties in the microphysical parameters $\epsilon_{e}$, $\epsilon_{T}$, and $\epsilon_{B}$. As discussed in §2 and above, we choose $\epsilon_{e}=0.01$, $\epsilon_{T}=0.4$, and $\epsilon_{B}=0.1$ as fiducial values for these parameters. Particle-in-cell (PIC) simulations generally support the idea that non-thermal particle acceleration occurs in collisionless shocks, however the efficiency of such acceleration (i.e. $\epsilon_{e}$) is subject to some uncertainty given numerical limitations on the run-time of these simulations. Some general trends are nevertheless noticeable and worth mentioning.
First, the acceleration efficiency may depend on the strength and orientation of any pre-existing upstream magnetic field (to be differentiated from the downstream shock-amplified field). In the ultra-relativistic $\Gamma\gg 1$ regime, magnetized shocks (where the ratio of magnetic to kinetic flux crossing the shock front is $\gtrsim 1$) inhibit non-thermal particle acceleration and cause $\epsilon_{e}\to 0$ so long as the upstream magnetic field is oriented at an angle $\gtrsim 0.6\,{\rm rad}/\Gamma\ll 1$ with respect to the shock normal (quasi-perpendicular shock; Sironi & Spitkovsky 2009). In the unmagnetized (or parallel shock) relativistic case, PIC simulations suggest that the non-thermal electron acceleration efficiency may be as high as $\epsilon_{e}\sim 0.1$ (e.g., Sironi et al. 2013). This is consistent with $\epsilon_{e}$ values that have been recently inferred for GRB radio afterglows (Beniamini & van der Horst 2017; Duncan et al. 2023; although note that these results depend on modeling which neglects thermal electrons). Lower velocity, non-relativistic, shocks have also been extensively studied using PIC simulations. In this case the efficiency of non-thermal particle acceleration depends more sensitively on the orientation of any pre-existing upstream magnetic field. However it is generally found that for non-relativistic shocks $\epsilon_{e}$ is smaller than for relativistic shocks. This is broadly consistent with findings regarding the cosmic-ray electron-to-proton ratio which imply that $\epsilon_{e}\ll\epsilon_{p}\sim 0.1$ (e.g., Morlino & Caprioli, 2012).
Our present choice of fiducial $\epsilon_{e}=10^{-2}$ is more in line with expectations for non-relativistic or mildly-relativistic shocks. If $\epsilon_{e}$ is larger in the relativistic regime then this may lead to a situation where $\epsilon_{e}\sim\epsilon_{T}$ for ultra-relativistic shocks. In this scenario, the effect of thermal electrons would be less significant because there would be little dynamic range (in frequency) where thermal electrons dominate over power-law ones. For example, in the optically-thin (relativistic) SED shown in the right-hand panel of Figure 1, the transition to power-law-dominated emission (where the SED follows $L_{\nu}\propto nu^{-(p-1)/2}$) would happen much closer to $\nu_{\rm pk}=\nu_{\Theta}$ if $\epsilon_{e}\sim\epsilon_{T}$. This would inhibit the ‘thermal bump’ and make the SED and light-curve look very-much like standard power-law models. We speculate that this may be one reason why there is, currently, no strong evidence supporting the presence of thermal electrons in GRB afterglows (although we emphasize that this deserves more attention in the literature).
The findings from PIC simulations that $\epsilon_{e}$ may depend on the shock velocity, increasing across the transition from non-relativistic to ultra-relativistic shocks, has important implications. Introducing a velocity-dependent $\epsilon_{e}(\Gamma\beta)$ would affect our results in various ways, as was first pointed out in Margalit & Quataert (2021). This could be incorporated into our current framework, but is beyond the scope of this paper—particularly given the significant uncertainties and poor quantitative understanding of the relation $\epsilon_{e}(\Gamma\beta)$. It may be useful to revisit this in the future.
The detailed mechanisms and efficiency of electron-ion equilibration (and therefore the value of $\epsilon_{T}$) are also not fully understood at present. Coulomb collisions are generally ineffective given their associated long timescales. Nevertheless, PIC simulations of shocks generically show that electrons gain energy and establish quasi-equilibrium with ions, even though the detailed plasma processes responsible for mediating this energy transfer are still debated. Examining the limiting case where electrons do not exchange energy with ions at all places a strict lower limit of $\epsilon_{T}>\mu_{e}m_{e}/\mu m_{p}\simeq 10^{-3}$. PIC simulations of relativistic shocks suggest instead that $\epsilon_{T}\sim O(1)$ (e.g., Sironi et al. 2013; Groselj et al. 2024; Vanthieghem et al. 2024). In the non-relativistic regime there is evidence that $\epsilon_{T}$ may be lower, and in particular data from SN remnants and from the solar wind suggest a strong dependence on the shock Mach number (e.g., Sironi et al., 2013; Tran & Sironi, 2020; Raymond et al., 2023).
Given the current lack of a firm understanding of $\epsilon_{T}$ and its potential velocity dependence, we have adopted a constant $\epsilon_{T}=0.4$ that is consistent with the naive scenario where electrons equilibrate with ions. This is also in line with recent work by Vanthieghem et al. (2024) which uses PIC simulations to study electron-ion equilibration and finds that $\epsilon_{T}\approx 0.3-0.4$ over a wide range of shock velocities $0.075\lesssim\left(\Gamma\beta\right)_{\rm sh}\lesssim 3\times 10^{4}$.777 Vanthieghem et al. (2024) define a parameter $\alpha$ such that $kT_{e}\equiv\alpha\left(\Gamma_{\rm sh}-1\right)\mu m_{p}c^{2}$. They find that $\alpha\approx 0.1$ nearly independent of the shock velocity. This parameter is related to our $\epsilon_{T}$ via $\epsilon_{T}=8\alpha/3$ in the non-relativistic regime, and $\epsilon_{T}=3\sqrt{2}\alpha$ for ultra-relativistic velocities. Our assumption that $\epsilon_{T}=0.4$ can be relaxed by inserting lower values of $\epsilon_{T}$ in any of the equations in this paper, and this will generally have a significant impact on many of the results. A corollary of the uncertainty regarding $\epsilon_{T}$ is that our work makes $\epsilon_{T}$-dependent predictions that may help probe these microphysics and constrain $\epsilon_{T}$.
@@ 7 Summary
  • For shocks governed by SSA: thermal electrons dominate the emissivity and absorption at frequency $\nu_{\rm pk}$ when $L_{\nu_{\rm pk}}>L_{\rm th}$ (Equation 8; solid grey curve in Figure 2). For luminous events that fall within this regime, the standard Chevalier (1998) relations (4.1,4.1) are no longer applicable. Instead, the physical properties of the shock should be estimated from Equations (4.2,4.2; see also Appendix A) or using the numerical procedure outlined in Appendix B (implemented in code available at https://github.com/bmargalit/thermal-synchrotron-v2; see also Margalit 2024).
@@
  • There is a maximum allowed peak luminosity $L_{\nu_{\rm pk}}$, such that $L_{\nu_{\rm pk}}\leq L_{\rm crit}$ (Equation C7; see also Equation 5.1 and Figure 2). This is a consequence of the transition between optically-thin and optically-thick emission at the peak of the SED, and is a generic feature of synchrotron shock models.888 The expression for $L_{\rm crit}$ is determined by properties of the thermal electron population (§5; Appendix C). However, a qualitatively similar upper limit would apply also in the (unphysical) case where thermal electrons are neglected. This gives $L_{\rm crit}\sim 2\times 10^{33}\,{\rm erg\,s}^{-1}\,{\rm Hz}^{-1}\,\epsilon_{e,-2}^{-2}\epsilon_{B,-1}^{-23/55}(\nu_{\rm pk}\bar{t}/5{\rm GHz}\times 100{\rm d})^{133/55}$. While the specific luminosity at low frequencies $\nu\ll\nu_{\rm pk}$ can exceed this limit (that is, $L_{\nu}>L_{\rm crit}$ is possible) when peak emission is optically-thin, the peak luminosity should always satisfy $L_{\nu_{\rm pk}}\leq L_{\rm crit}$. It is interesting to note from Figure 2 that no observed event falls above this limit (red curve), despite clear observational biases towards detecting more luminous events. We argue that the limit $L_{\rm crit}$ that we have derived here provides a natural explanation to this fact.
@@
The code developed in this analysis is publicly available (Margalit, 2024) and may be particularly useful when the analytic approximations derived in the main text are not sufficiently accurate. This code can be imported as a python module and used to infer physical properties of observed transients, for example by fitting model-generated flux densities to real data. It may also be useful for generating plots similar to Figures (2,5).
B.M. thanks Nayana A. J. for sharing a compilation of $\nu_{\rm pk}$ and $L_{\nu_{\rm pk}}$ for SNe which provided a basis for the sample of events shown in Figures 2,5. This research benefited from interactions that were funded by the Gordon and Betty Moore Foundation through Grant GBMF5076. (Harris et al., 2020), scipy (Virtanen et al., 2020), matplotlib (Hunter, 2007), thermalsyn (Margalit, 2024).
@@ Appendix B General Solution
are functions of the shock velocity and effective mass-loss rate. In the above, $f(\Theta)$ and $g(\Theta)$ (not to be confused with the volume filling factor $f$) are functions of the normalized electron temperature and are given by Equations (5,8) of Margalit & Quataert (2021).999 Note that there was a typo in the expression for $g(\Theta)$ presented in Equation (8) of Margalit & Quataert (2021). The correct expression is $g(\Theta)=\frac{(p-1)\gamma_{m}}{(p-1)(\gamma_{m}-1)-p+2}\left(\frac{\gamma_{m}}{3\Theta}\right)^{p-1}$. The electron temperature is itself a function of the shock velocity, and is determined by Equation (2) of Margalit & Quataert (2021) with $\Theta_{0}=\epsilon_{T}(\mu m_{p}/\mu_{e}m_{e})(\Gamma-1)$. In the limit $\Theta\gtrsim 1$ which is relevant for high-velocity shocks, $f(\Theta)\approx g(\Theta)\approx 1$. The coefficients $C_{j}(p)$ and $C_{\alpha}(p)$ are given by Equations (15,17) of Margalit & Quataert (2021). For $p=3$, these attain the values $C_{j}=2/3$ and $C_{\alpha}\simeq 28.6$. Finally, the function $I^{\prime}(x)$ is given by Equation (13) of Margalit & Quataert (2021), and is derived in Mahadevan et al. (1996).